Two number cubes are rolled. What is the probability of rolling either a total greater than 9 or a multiple of 5?
Answer by greenestamps(13200)



P(5 or 10 or 11 or 12) = P(5)+P(10)+P(11)+P(12) = 4/36+3/36+2/36+1/36 = 10/36 = 5/18